# Corrigendum for Han’s Corrigendum

### v1

@@ 4 Other flaws in [19]
  • Suppose $b=c_{j}$ such that $j\neq 0\mod n$. We have
    $p^{-1}(N(c_{j}))=\bigcup_{i=0}^{\infty}[i\cdot j-1,i\cdot n+1]_{{\mathbb{Z}}}=\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i}),$
    where $i\mod n=j$ and $p^{-1}(c_{j})=\{e_{i}\mid i\in{\mathbb{N}}^{*}\}$.
@@ 5 Further remarks
Authors are responsible for their writings, but referees who approve publication of mathematical trash also bear some responsibility. The point is not that a Han submission should be automatically rejected - he deserves credit for useful ideas such as the introduction of covering maps and the shortest-path metric into digital topology. Rather, his submissions should be reviewed with the greatest care, by reviewers with good knowledge of the relevant literature, who are not afraid to recommend either rejection or major revision.

### v2

@@ 4 Other flaws in [19]
  • Suppose $b=c_{j}$ such that $j>0$. We have
    $p^{-1}(N(c_{j}))=\bigcup_{i=0}^{\infty}[j+i\cdot n-1,j+i\cdot n+1]_{{\mathbb{Z}}}=\bigcup_{i\in{\mathbb{N}}^{*}}N(e_{i}),$
    where $e_{i}=j+i\cdot n$ and $p^{-1}(c_{j})=\{e_{i}\mid i\in{\mathbb{N}}^{*}\}$.
@@ 5 Further remarks
Authors are responsible for their writings, but referees who approve publication of mathematical trash also bear some responsibility. The point is not that a Han submission should be automatically rejected - he has introduced useful ideas into digital topology, such as covering maps and the shortest-path metric. Rather, his submissions should be reviewed with the greatest care, by reviewers with good knowledge of the relevant literature, who are not afraid to recommend either rejection or major revision.
6 Acknowledgment
We are grateful to the anonymous referees for their suggestions and corrections.